Discrete Math

Find d = gcd(131, 122). Find integers s and t such that d = 131 . s + 122-t
Answer #1

Now 131= 1 À 122 + 9 122 = 1389 +5 9 = 1X5+4 0 5= 1X4+1 4= 4x170 goeid 11311122) = 1 2 HonceNow I= 1 X 5 - 144 5 -|(9-5) . = -9 + 2x5 1 =-9+2 (122713.89). I = -9 + 2x122-2689 I. -9749 + 2X122 1. - 27(131-122) + 2x122